### v2

# Impact of boiling liquid droplets: Vapor entrapment suppression

Abstract
There hardly is a fluid mechanics phenomenon attracting more attention than the impact of a droplet, due to its undeniable beauty, many applications and the numerous challenges it poses. One of the crucial factors turns out to be the cushioning effect of the gas surrounding the droplet. This fact, together with the observation that almost all of the relevant literature was done in air, triggers the question what would happen when the liquid was a boiling liquid, i.e., a liquid in thermal equilibrium with its own vapor, as is the case during transport of cryogenic liquids such as liquid hydrogen. To investigate precisely this question, we experimentally generate droplets in thermodynamical equilibrium with their own vapor, even before impact, such that minute energy exchanges of the droplet with its surroundings can trigger phase change. Using a frustrated total internal reflection (TIR) setup, we make the exciting observation that depending on the impact speed and vapor conditions, the entrapment of vapor can be completely suppressed under boiling liquid conditions. We create a simplified model based on scaling arguments and perform numerical simulations considering both the compressible and condensable properties of the vapor layer that are in very good agreement with our experimental findings. Our results can be of great consequence to the pressures exerted during droplet impact and on an industrial scale may help better understand the loads experienced during sloshing wave impact inside cryogenic liquid containers.
Droplet impact has long remained a subject of interest to fluid mechanics for its beauty, applications and numerous challenges [34, 11]. In fact, when it comes to studying impacting droplets, several factors can affect the sequence of events which will unfold, such as the substrate’s roughness [13, 6], or whether the liquid wets it or not [25] and even the ambient pressure, which can suppress the characteristic splash of high velocity droplet impacts [33]. After the discovered relevance of the ambient pressure, of course related to the gas surrounding the droplet, recent attention has been brought to the previously neglected interaction of the droplet with the gas layer between the liquid and solid which has led to many relevant findings. Roughly speaking, when a falling droplet gets close enough to a surface, pressure builds up in the gas layer below the droplet’s bottom surface and the gas underneath is squeezed out radially. This leads to the formation of a dimple shape along the droplet’s lower surface [19] which will have relevant consequences later in the impact. One of these is that the very early interaction with the gas layer can later lead to splashing [20, 26], another is that the dimple shape can become an entrapped gas bubble after touchdown [1] or it can lead to jetting, either upwards [16] or downwards [18], which in turn may also change the dynamic impact force on the surface [35]. With all this in mind, it can be easily concluded that the gas layer interaction still holds great importance in the research of impacting droplets and its many applications.
In this work we replace gas by vapor, by focusing on the impact of boiling liquid droplets, that is, droplets that are at their boiling point, in thermal equilibrium with the vapor that surrounds them. More specifically, we want to focus on the entrapment of vapor inside the dimple formed below the droplet due to pressure accumulation. A related subject is that of droplets impacting in (or close to) the well-known Leidenfrost regime [24], where due to evaporation a droplet can levitate over a heated solid upon impact. These have been performed in atmospheric [30, 17, 8, 2, 7, 28] and reduced pressure conditions [31]. Unlike this previous work, we do not use large temperature differences to trigger rapid phase change, since the liquid will be in thermodynamical equilibrium with its own vapor, even before impact. When in such a state, minute energy exchanges of the droplet with its surroundings can trigger phase change [4]. Moreover, unlike the Leidenfrost impact work where phase change is limited to the evaporation of the liquid due to high substrate temperatures, in this boiling liquid state the condensation of vapor due to pressure locally rising above the vapor pressure of the liquid is also possible. We will show that the latter may lead to dramatic effects.
This study is curiosity driven, but with a specific application in mind: During the maritime transport of cryogenic liquids, such as liquid natural gas (LNG) or liquid hydrogen (LH2), a liquefied gas is loaded onto a specialized carrier and hauled for long periods of time. Naturally, such a liquid is what we have defined above to be a boiling liquid: a liquid at its boiling point in thermal equilibrium with its own vapor. One of the main concerns during transport is to insulate the main container, in order to keep a low boil-off rate of the liquid to avoid losses of product. In this engineering challenge, the trade off for good insulation is loss of structural integrity, turning loads due to liquid impact inside of the container into a major concern [3]. While LNG carrier design has now long been known to work in a safe manner [32], the loads within it are not well understood, which is especially true for possible effects of phase change. This work aims to help understand such loads, by reducing the problem to the more fundamental one of droplet impact in equilibrium conditions that reflect those within cryogenic containers. Further research can help the transition to more sustainable fuels, such as LH2, which will pose further, more complex challenges.
Experimental setup
In order to generate droplets in equilibrium with their own vapor and record their impact onto a substrate, we have developed a highly versatile setup, using HFE-7000 as a working fluid, which has an appreciable vapor pressure ($P_{V,0}=0.57$ bar) at room temperature. Here we provide a general description of the setup (see Fig. 1), while a more detailed description can be found in the Supporting Information (SI).
Figure 1: Sketch of the experimental setup with as main elements the closed, temperature-controlled chamber containing HFE-7000 liquid (Res1) in equilibrium with its vapor, the droplet generated from the second reservoir (Res2), and the prism (P) used for frustrated total internal reflection imaging (TIR), the vacuum pump (VP), and the two high-speed cameras (CAM1,2) used for imaging.
First, a saturated liquid-vapor environment at temperature $T_{0}\approx 22$ C is created by allowing HFE-7000 to evaporate inside a hermetically sealed, vacated, and temperature-controlled chamber. Enough liquid is allowed into a reservoir (Res1) to ensure the chamber remains saturated with vapor during the entire experiment. Additional HFE-7000 liquid is allowed into a smaller reservoir (Res2), located higher in the setup, connected to a needle (N) with its tip slightly below the liquid level, allowing droplets to fall down and impact a sapphire prism (P). The height between needle and substrate can be varied to obtain an impact velocity $U_{0}$ between $0.2$ m/s and $1.5$ m/s. The substrate temperature $T_{s}$ is controlled between $T_{0}$ and $40$ C.
Side views of the impact are recorded at 16,000 fps and are used to measure the impact speed $U_{0}$ and initial radius $R_{0}$ of the droplets. Bottom views are captured at frame rates between $96,000$ and $200,000$ fps by illuminating the top surface of the prism internally with a collimated monochromatic light source (CLS2) from below, such that it totally reflects. If a droplet comes within the evanescent wave distance from the surface, then light is partly, and upon contact fully transmitted into the droplet, which results in dark regions on the otherwise bright surface. This technique is known as frustrated total internal reflection (TIR) and has been extensively used in the past [27, 14, 2].
Results and discussion
We have performed an extensive experimental campaign, divided in two parts. For the first part we have performed experiments where boiling droplets, falling at different velocities, impact a slightly heated surface. The latter is done to avoid the formation of a liquid film on top of the substrate that would interfere with the impact. For simplicity, we will refer to the experiments in this first part as (near-)isothermal experiments.
Near isothermal impact of boiling droplets
Fig. 2A shows TIR snapshots from typical near-isothermal experiments for droplets impacting a substrate at different impact speeds. The frames are a TIR sequence where wet (dark) and dry (light) zones are well distinguished. For $U_{0}=0.38$ m/s, in the first frame, a dark ring appears over the background, corresponding to the surface of the droplet making contact with the substrate. The light part within the ring is the dimple which has formed below the droplet, indicating entrapped vapor. As the droplet further moves down, the ring expands radially, both inwards and outwards.
This matches the chain of events observed for experiments conducted when the surrounding fluid is a non-condensable gas, as described in several other works [29, 1, 15, 18, 35, 12]. So we stress that a boiling liquid also entraps the vapor surrounding it, at least at this small impact velocity. However, increasing the impact velocity can lead to a radical change in behavior, which we can observe for the larger impact velocities in Fig. 2A.
[…]
Figure 2: (A) TIR image sequence (after background subtraction) of a series of HFE-7000 droplets with radius $R_{0}\approx 0.6$ $m$ in a vapor atmosphere at $T_{0}\approx 22$ C; the droplets impact a sapphire substrate heated to a temperature difference $\Delta T\lesssim 2$ C with impact velocities $U_{0}=0.24$, $0.38$, $0.51$, $0.74$ m/s, Fig. S2 and movies S1 to S5 in the SI. At low impact velocity a vapor bubble is entrapped, but at high impact velocities this entrapment is suppressed. (B) TIR image sequence (after background subtraction) of a series of droplets with radius $R_{0}\approx 0.6$ $m$ in a vapor atmosphere at $T_{0}\approx 22$ C impacting with velocity $U_{0}\approx 0.7$ m/s onto a sapphire substrate heated to an increasing temperature difference $\Delta T=2.4$, $6.4$, $11.3$, $17.6$ K (with respect to $T_{0}$), Movies S6 to S10 in the SI. At this impact velocity, the entrapment of vapor is suppressed for small $\Delta T$, but entrapment can be recovered by heating the substrate to a higher temperature. In all cases, time is measured with respect to the first TIR frame in which the droplet becomes visible.
In order to better quantify the transition to the suppression of vapor entrapment we have measured the initial contact radius $r_{0}$ of the entrapped region, which is an indication of the size of the entrapped vapor bubble, and the entrapment time $t^{*}$, the time interval that the bubble is visible in the TIR images. To do this, we obtain the inner and outer radii $r_{b}$ and $r_{d}$ of the ring shape during the initial contact and track their time evolution, as exemplified in Fig. 3A. The outer radius $r_{b}$ is the spreading contact line of the wetting region, which is known to expand as $r_{b}\sim\sqrt{t}$ [26]. This behavior is fitted in Fig. 3A and shows excellent agreement to the data. The retracting contact line (inner radius) is approximated using a linear fit. We estimate the initial contact time and radius $r_{0}$ by determining the intersection point of the fitted curves. Additionally, we also find the zero crossing of the linear fit to the retracting inner contact line to obtain the vapor entrapment time $t^{*}$, which we measure with respect to the initial contact time (Fig. 3A).
The plots in Figs. 3B and 3C present the measured initial entrapment radius $r_{0}$ and entrapment time $t^{*}$, respectively, for the isothermal experiments. Each point in the plot is the mean value found for at least three repetitions and the vertical bars show the total error with respect to the mean. Both the initial entrapment radius and the entrapment time decrease monotonically with impact velocity, as was already clearly visible from Fig. 2A. The initial contact radius is smaller than what has been previously reported for droplets surrounded by a non-condensable gas. It also presents a sudden decrease around $U_{0}\approx 0.4$ m/s and is only measurable for impact velocities smaller than $0.6$ m/s. Impacts past this velocity result in no entrapment, such as for the largest velocity in Fig. 2A. Most notably, at the threshold velocity around $0.6$ m/s, the entrapment time goes to a small asymptotic value that is consistent with zero, whereas the entrapment radius remains finite.
Figure 3: (A) Azimuthally averaged inner ($r_{b}$, blue points) and outer ($r_{d}$, orange points) radius evolution for an experiment at $U_{0}=0.38$ m/s, with images available in the SI, Fig. S1B. The orange line is obtained from a linear fit to $r_{d}^{2}$ plotted against $t$ and the blue line is a linear fit for $r_{b}$. The intersection of the two curves provides the initial radius $r_{0}$ and also the initial impact time. The entrapment time $t^{*}$ is then measured from the inner radius fit as the time it takes for $r_{b}(t)$ to go from $r_{0}$ to 0. (B) Initial entrapment radius $r_{0}$ as a function of the impact velocity $U_{0}$ for a set of experiments with a small temperature difference of the impact surface with respect to the ambient temperature ($\Delta T=1.77\pm 0.49$ K). The points represent the mean of a minimum of three repetitions and the bars display the total error. (C) entrapment time $t^{*}$ as a function of the impact velocity $U_{0}$, for the same set of experiments.
A simplified one-dimensional model
In the next few paragraphs, through a simplified model, we find a quantitative explanation for the observed suppression of vapor bubble entrapment for increasing impact velocity. We begin considering an impacting droplet surrounded by a non-condensable gas. The pressure difference under such droplet is described by the lubrication flow of the gas as the droplet nears the surface [21, 1]. As argued in that work, we can estimate the pressure build up below the droplet from the axisymmetric Stokes equation in the lubrication layer as
$\frac{\Delta P}{\ell}\sim\mu_{g}\frac{U_{g}}{h^{2}}\quad\Rightarrow\quad\Delta P\sim\mu_{g}\frac{R_{0}}{h^{2}}U_{0}\,\,$, (1)
where $\mu_{g}$ is the dynamic viscosity of the gas, $h$ is the thickness and $\ell\sim\sqrt{R_{0}h}$ the radial extent of the lubrication layer, where the latter is estimated geometrically from the approach of a spherical droplet to a planar substrate, and the gas velocity $U_{g}$ is related to the impact velocity by continuity, $U_{g}h\sim U_{0}\ell$. By equating this $\Delta P$ with the pressure needed to decelerate the liquid, namely $\rho_{L}(\partial\phi/\partial t)\sim\rho_{L}U_{0}^{2}\ell/h$ (with $\phi$ the flow potential), one obtains a scaling relation for the dimple height $h_{d}$ in terms of the Stokes number $\mathit{St}\equiv\rho_{L}U_{0}R_{0}/\mu_{g}$
$\frac{h_{d}}{R_{0}}\sim\left(\frac{\rho_{L}U_{0}R_{0}}{\mu_{g}}\right)^{-2/3}=\mathit{St}^{-2/3}\,\,$, (2)
where $\rho_{L}$ is the density of the liquid.
For our case, however, the surrounding phase is vapor and this means that an increase in pressure can trigger condensation: As soon as the pressure in the vapor layer rises, the vapor will try to condense to return to the equilibrium state. For simplicity, let’s now consider a one-dimensional system composed of a solid substrate, an incoming liquid surface and an intermediate vapor layer. During pressure build up, vapor will try to condense and produce latent heat. Since simultaneously the temperature in the vapor rises, a thermal boundary layer will develop in the liquid, which will allow for transport of the latent heat into the liquid. Now, writing down the mass-energy balance at the liquid-vapor interface [23], an approximation we also used in [4, 5]
$L\frac{dm_{V}}{dt}=-S\,k_{L}\left.\frac{\partial T}{\partial r}\right|_{\text{int}}\approx S\,k_{L}\frac{\Delta T_{V}}{\delta_{th}}\,\,$, (3)
where $L$ is the latent heat, $dm_{V}/dt$ the rate of change of vapor mass due to condensation, $S$ the surface area of the liquid-vapor interface, $k_{L}$ the thermal conductivity of the liquid, and $\partial T/\partial r|_{\text{int}}$ the temperature gradient at the interface. Note that we have neglected the amount of liquid produced during condensation due to the large liquid to vapor density ratio, and the sensible heat of the vapor in light of the much larger latent heat. Subsequently, we estimate the temperature gradient as the ratio of $\Delta T_{V}$, the difference between the temperature $T_{V}$ of the vapor and the equilibrium temperature $T_{0}$ of the liquid far away from the interface, and $\delta_{th}$, the thermal boundary layer thickness in the liquid phase.
The condensation rate $dm_{V}/dt$ of the vapor can be interpreted as a condensation speed in the vapor phase by dividing by vapor density $\rho_{V}$ and surface area $S$, i.e., $U_{cond}=-(dm_{V}/dt)/(\rho_{V}S)$, which after substituting in the previous equation gives us
$U_{cond}\approx\frac{k_{L}}{L\rho_{V}}\frac{\Delta T_{V}}{\delta_{th}}\,\,.$ (4)
Assuming that the vapor bubble is in internal thermal equilibrium at all times (connected to neglecting the influence of sensible heat) the relation between vapor temperature and pressure is described by the Clausius-Clapeyron equation and the ideal gas law, for which one can write in linear approximation that
$\frac{\Delta T_{V}}{T_{0}}\approx\beta\frac{\Delta P_{V}}{P_{V,0}},\quad\text{with:}\quad\beta\equiv\frac{R_{s}T_{0}}{L}\,\,$, (5)
where $R_{s}$ is the specific gas constant of the vapor phase and the equilibrium state variables of the vapor are denoted with subscript 0: $P_{V,0}$, $\rho_{V,0}$, and $T_{0}=T_{V,0}$. Finally, the boundary layer thickness $\delta_{th}$ can be approximated as $\delta_{th}\sim\sqrt{\pi\alpha_{L}\Delta t}$, with $\alpha_{L}$ the thermal diffusivity of the liquid and the time scale $\Delta t\approx\epsilon\frac{R_{0}}{U_{0}}$, where $\epsilon$ should be approximated from evaluating the time that the pressure rise in the lubrication layer starts being appreciable. This leads to $\epsilon\approx 0.02$, corresponding to $h\approx 0.02R_{0}$, which is justified in the SI and is both consistent with a similar estimate in [8] and with BI simulations (see next Section). Substituting everything into the expression for the condensation speed, and dividing by $U_{0}$ we obtain
$\frac{U_{cond}}{U_{0}}\approx\frac{\beta^{2}}{\sqrt{\pi\epsilon}}\frac{\rho_{L}}{\rho_{V,0}}\frac{c_{p,L}}{R_{s}}\sqrt{\frac{\alpha_{L}}{R_{0}U_{0}}}\frac{\Delta P_{V}}{P_{V,0}}\,\,$, (6)
where we used $k_{L}=\alpha_{L}\rho_{L}c_{p,L}$, with $\rho_{L}$ and $c_{p,L}$ the density and specific heat of the liquid.
Now the main argument to connect the non-condensable gas derivation and the condensation speed derived above is the following: If at the moment that in a non-condensable gas situation the dimple would be formed, i.e., if (1) and (2) would both hold, such that
$\Delta P_{V}(\approx\Delta P_{g})\sim\mu_{V}\frac{R_{0}}{h_{d}^{2}}U_{0}\sim\frac{\mu_{V}U_{0}}{R_{0}}\mathit{St}^{4/3}\,\,$, (7)
the condensation speed $U_{cond}$ is larger than the impact velocity $U_{0}$, then the vapor layer would condense faster than the impact speed and no vapor bubble will be entrapped. Conversely, if $U_{cond}<U_{0}$, we will see vapor bubble entrapment. Inserting (7) into (6) leads to
$\frac{U_{cond}}{U_{0}}\sim\frac{\beta^{2}}{\sqrt{\pi\epsilon}}\frac{\rho_{L}}{\rho_{V,0}}\frac{c_{p,L}}{R_{s}}\frac{\mathit{Eu}\,\mathit{St}^{1/3}}{\mathit{Pe}^{1/2}}\,\,$, (8)
where $\mathit{Pe}=R_{0}U_{0}/\alpha_{L}$ and $\mathit{Eu}=\rho_{L}U_{0}^{2}/P_{V,0}$ are the Péclet and the Euler numbers, respectively. Note that we have replaced $\mu_{g}$ with $\mu_{V}$ to clearly express that the dynamic viscosity of the vapor is intended.
Now, the threshold condition $U_{cond}/U_{0}=1$ can be computed by turning the approximate result (8) into an equation using a multiplicative numerical constant $K_{0}$ in the right hand side and subsequently inserting literature values for the properties of HFE-7000 at $T_{0}=22$ C (see SI). For the reasonable value $K_{0}=1.6$ this results in $U_{0}\approx 0.52$ m/s, in fair agreement with our experimental findings. This is quite remarkable noting the number of approximations that enter the above derivation.
Numerical simulations
To further support our findings, we have performed a series of simulations using an in-house boundary integral (BI) method, based on the code first described in [22]. The method has proven to be reliable for solving potential free surface flows and has been previously adapted to simulate drop impact [1]. For the numerical simulations, the flow inside the droplet is assumed to be a potential flow and the droplet shape to be axisymmetrical. This system is solved with the BI method to yield the time evolution of the droplet contour, where it is coupled to a compressible and condensable lubrication flow in the vapor layer. This is in contrast to the work in [1], where for the non-condensable air layer it had been coupled to an incompressible lubrication flow, and to [10], where a different phase change model was used. More details of the simulation technique and the numerical setup can be found in the SI.
We performed numerical simulations for similar conditions as the experimental ones reported in Fig. 2A, starting the simulation from an (arbitrary) height $h_{0}$, large enough to not influence results. More specifically, in Fig. 4A we extract the shape of the lower portion of an impacting droplet from the simulations. For each case, subsequent profiles are shown in time intervals of $1.33\cdot 10^{-3}R_{0}/U_{0}$, such that, far from the substrate, successive profiles for different $U_{0}$ would be equidistant, and the last profile corresponds to just before the simulation has stopped due to overlapping with the substrate.
Figure 4: (A) Time evolution of the axisymmetric and normalized profile $h/R_{0}$ as a function of the normalized radial coordinate $r/R_{0}$ for four boundary integral simulations of a HFE-7000 droplet impacting at different speeds $U_{0}=0.2$, $0.4$, $0.6$, $0.8$ m/s. The horizontal black dashed lines at $h=455$ nm indicate the height at which a droplet would enter the evanescent wave of the light source, and would thus show up in a TIR experiment. Color indicates time progression from yellow to dark blue. (B) Time evolution of the center height $h_{c}$, normalized with the initial height $h_{0}$, for various impact velocities $U_{0}$ between $0.2$ m/s and $1.0$ m/s (blue to yellow). The straight dashed black line represents the time evolution of the position of the bottom of a moving sphere in the absence of vapor. Inset: Zoom into the blue dashed rectangular area. (C) The center touchdown time $\Delta t$ (blue pentagons), plotted as a function of the impact velocity $U_{0}$, where the continuous red curve is a hyperbolic fit to the data and the dashed red vertical line represents its asymptote. The red stars represent the entrapped vapor volume $V_{\text{e}}$ indicated in (A).
For a small impact velocity ($U_{0}=0.2$ m/s in Fig. 4A) a clear dimple is formed, which will entrap vapor upon touchdown: Starting from the first (upper) profile, which is still monotonically decreasing towards the center ($r/R_{0}=0$), one observes that the surface in the center is halted and only moves downward slowly, forming a dimple in the center, which subsequently will lead to the entrapment of a vapor bubble. After the last profile shown, the liquid will make contact with the substrate at the minima in the profile. For the next impact velocity, at $U_{0}=0.4$ m/s, we see a similar profile sequence as for $U_{0}=0.2$ m/s, with some subtle differences: The dimple formation happens closer to the surface, and towards the end the central part of the profile moves downward slightly faster.
At $U_{0}=0.6$ m/s, a qualitative difference appears in the dynamics of the bottom surface. From the profile sequence it becomes clear that also in this case the profile in the center is decelerated, but unlike what happened in the previous case, it continues to move downwards, and the bottom is nearly flat such that, even if some vapor is entrapped upon impact, it is likely to be very short lived: The subsequent build-up of pressure in the entrapped bubble will be able to quickly condense whatever vapor is left inside it. Finally, at the highest impact velocity ($U_{0}=0.8$ m/s in Fig. 4A) there is no sign of dimple formation at all: The droplet falls through the vapor layer as if no intermediate phase existed and hits the substrate with its center, without entrapment of a vapor bubble. Note that the spacing between profiles becomes larger with increasing speed, indicating considerably less deceleration for large $U_{0}$. All these findings are consistent with the experimental data presented in Fig. 2A. In the SI (Fig. S3) we compare the profiles of Fig. 4A to those of a non-condensable gas with the same properties.
Summarizing, in both experiment and simulation we find similar behavior: For small impact velocities a dimple is formed that leads to the entrapment of a vapor bubble, at intermediate velocities ($U_{0}\approx 0.5$ m/s) the dimple shape becomes a lot less prominent, consistent with the speed where we no longer observe entrapment in the experiments. For the highest impact velocity shown, neither a dimple shape, nor an entrapped vapor bubble can be discerned at all.
We now proceed to use the simulations to determine the transition impact velocity $U_{0,\text{thr}}$ from vapor entrapment to immediate touchdown of the liquid surface. To quantify $U_{0,\text{thr}}$, we determined the time difference $\Delta t$ between the moment of impact of the bottom of a sphere approaching the substrate in the absence of an intermediate phase and the impact time of the centre of the bottom surface of a simulated drop impact. To obtain the latter, we have tracked the center height $h_{c}$, which is the distance from the substrate to the bottom liquid surface at the center, i.e., at $r=0$, for simulations with impact velocities $0.2\leq U_{0}\leq 1$ m/s, as shown in Fig. 4B. As already mentioned, the simulations will stop a little before contact, such that we do not possess numerical data all the way to touchdown and linearly extrapolate the dimple height to determine $\Delta t$ as indicated in the inset of Fig. 4B.
[…]
Impacts on a moderately superheated surface
To complement the isothermal experiments, we have performed experiments with a significantly heated substrate, which in principle should counteract the condensation. This is reminiscent of the well-known Leidenfrost effect [24], where a droplet is made to hover over a strongly heated substrate levitated by its own vapor. The result is shown in Fig. 2B, for droplets impacting at $U_{0}\approx 0.7$ m/s. For superheat $\Delta T=T_{s}-T_{0}=2.4$ K, the entrapment is completely suppressed, as based on our isothermal experiments would be expected for this impact velocity. However, as soon as we increase the substrate temperature the droplet will start to entrap vapor again. For four degrees higher substrate temperature ($\Delta T=6.4$ K) the lifetime of the vapor bubble is extremely brief (less than 10 $\mu s$), the initial entrapment radius small, and moreover only visible in the first frame. At $\Delta T=11.3$ K, the entrapped vapor bubble remains visible in the first two frames and further incrementing $T_{s}$ to $\Delta T=17.6$ K we fully recover vapor entrapment: An axisymmetric wetted ring is found in the first recorded frame and the inner and outer contact lines expand as was previously observed in the isothermal low impact velocity experiments, where the entrapped vapor bubble remains visible up to the last frame shown in the figure.
Figure 5: Experimental phase diagram tracing the boundary between droplet impacts with (colored circles) and without (crosses) vapor entrapment. On the vertical axis the impact velocity $U_{0}$ is plotted, whereas on the horizontal axis we find the temperature difference $\Delta T$ of that of the heated substrate and the equilibrium temperature of the surrounding vapor. The color of the circles denotes the bubble lifetime $t^{*}$, as indicated by the color bar. The blue curve indicates the phase boundary predicted from the model.
In summary, the process by which entrapment of a vapor bubble below a droplet impacting at higher velocities is suppressed, can be reversed by heating the substrate, frustrating the condensation that occurs at the vapor-liquid interface. The phase diagram in Fig. 5 traces the transition from vapor bubble entrapment to its suppression in the parameter space formed by the substrate superheat $\Delta T$ and the droplet impact velocity $U_{0}$. Black crosses denote impacts without entrapment and circles indicate that a vapor bubble was entrapped, where their color represents the lifetime $t^{*}$ of the vapor bubble, changing from dark blue to bright yellow for increasing lifetime. For nearly isothermal impacts, no entrapment is recorded above $U_{0,\text{thr}}\approx 0.6$ m/s, in good agreement with the numerical simulations we conducted. When the superheat $\Delta T$ is increasing, this transition point keeps shifting to larger velocities. Notably, the entrapment lifetime along the transition line remains small, whereas bubbles corresponding to points further away from the transition line live longer.
Inspired by the analysis of the ordinary [24] and dynamic [8] Leidenfrost phenomenon, the superheat $\Delta T$ can be included by adding a parasitic linear heat conduction term through the vapor layer, $Sk_{V}(T_{s}-T_{V})/h\approx Sk_{V}\Delta T/h$ to the left hand side of (3), where $h=h_{d}$ is again estimated from the dimple condition (2). Using a second multiplicative numerical constant $Q_{0}$ in front of the parasitic heat conduction term, the condition $U_{cond}/U_{0}=1$ for $Q_{0}=1.35$ leads to the blue line in Fig. 5, which nicely demarcates the regions with and without vapor bubble entrapment in Fig. 5. Details of this calculation can be found in the SI.
Note that the transition from vapor bubble entrapment to no entrapment is gradual (second order) rather than abrupt (first order). This, in general, makes it more difficult to find the phase boundary in an experiment. In addition, the divergence observed in the touchdown time $\Delta t$ observed in the numerical simulation is an artificial one since condensation in principle continues and eventually the vapor bubble would disappear also in the simulation, which is why the dimple height bends back towards zero for the smallest impact speeds in Fig. 4B, leading to large but finite touchdown times. In the experiment there is the added effect (at least for small or zero superheat) that HFE-7000 is observed to be very wetting, which will lead to entrainment of the vapor bubble into the inside of the droplet.
Conclusions
In conclusion, we have performed impact experiments of droplets under boiling liquid conditions, i.e., of a liquid that is in thermal equilibrium with its own vapor. Whereas, for substrate temperatures close to that of the surroundings, at low impact speeds a vapor bubble is entrapped –similar to the entrapment of an air bubble for droplet impact under atmospheric conditions–, we observe that vapor bubble entrapment is suppressed completely at sufficiently high impact speeds. Heating the substrate, we can counteract this effect: If we start with a sufficiently high impact speed for which no vapor bubble is entrapped, we can by progressively heating the substrate reach a situation where vapor entrapment reappears.
We explain this behavior from condensation at the vapor-liquid interface: If the pressure build-up below the droplet is sufficiently high to cause the vapor to condense, at higher impact velocities we arrive in a situation where condensation becomes fast enough to consume the entire intermediate vapor layer during impact, and the droplet hits the surface without the cushioning action of the gas layer that is present during impact under atmospheric conditions.
We create an approximate model based on scaling arguments to explain our findings, which is in good agreement with our experiments. The argument starts from the same mass-energy balance approximation at the liquid-vapor interface we also used successfully in explaining the strongly increased pressures observed during wave [4] and disc [5] impact in a boiling liquid, which provides additional confidence in the correctness of this approach.
In addition, we devised a numerical simulation consisting of a boundary integral method for the impacting droplet, coupled to a compressible lubrication model including phase change for the vapor layer that provides insight into the vapor dynamics below the impacting droplet. Since gas compressibility effects in droplet impact are important and constitute an active research area [9, 28], it is good to stress that we are in an impact regime where compressibility effects are expected to be minor [21], as discussed in the SI. However, phase change (condensation) can only be reliably incorporated into a compressible lubrication model.
Our research is of consequence for the transport of cryogenic liquids such as LNG and LH2, where the suppression of vapor bubble entrapment is likely to go hand in hand with increased impact pressures. The models introduced in this work may be translated to the conditions in those applications to verify under which circumstances phase change may become a non-negligible effect and needs to be included into the engineering of cryogenic liquid carriers on an industrial scale. More specifically, using our theoretical results in estimating when the rate of condensation exceeds the impact speed of a sloshing wave crest, will make it possible to identify and avoid situations in which the protective cushioning effect of the intermediate vapor phase is cancelled by condensation. The latter will increase the impact load, potentially threatening the structural integrity of the transporting containers.